A gourmet coffee shop in downtown San Francisco is open 200 days a year and sells an average of 75 pounds of Kona coffee beans a day.
(Demand can be assumed to be distributed normally, with a standard deviation of 15 pounds per day.)
After ordering (Fixed cost=$16 per order), beans are always shipped from Hawaii within exactly four days.
Per-pound annual holding costs for the beans are $3.
a) What is the EOQ for Kona Coffee Beans?
b) What are the total annual holding costs of stock for Kona coffee beans?
c) What are the total annual ordering costs for Kona coffee beans?
d) Assume the management has specified that no more than a 1% risk during stockout is acceptable. What should the reorder point be?
e) What is the safety stock needed to attain a 1% risk of stockout during lead time?
f) What is the annual holding cost of maintaining the level of safety stock needed to support a 1% risk?
g) If management specified that a 2% risk of stockout during lead time is acceptable, would the safety stock holding costs decrease or increase?
Answer #1

EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST)

ANNUAL HOLDING COST = (EOQ / 2) * HOLDING COST

ANNUAL ORDERING COST = (DEMAND / EOQ) * ORDERING COST

A. EOQ = SQRT(2 * 15000 * 16 / 3) = 400

B. ANNUAL HOLDING COST = (400 / 2) * 3 = 600

C. ANNUAL ORDERING COST = (15000 / 400) * 16 = 600

Do give a positive rating if you think this answer helped